OpenStudy (anonymous):

Raja was a dentist and needed furniture for his surgery.He bought $6000 worth of furniture from david jones on hire purchased by paying 10% deposit and $540 a month for 1 year.How much extra did he have to pay?

OpenStudy (babalooo):

Something sounds funny with the david jones on hire purchased part... hmm...

OpenStudy (babalooo):

Hmmm.. can you double check if you typed the question correctly? If i'm understand it correctly...Raja purchased a total $6000 worth of furniture but cannot pay it all at once so he's paying it in parts. Okay... Assuming the the 10% deposit is a one time payment, first calculate what the deposit amount is. 10% of 6000 is a tenth of 6000. What is that? Now In addition to that deposit, he pays $540 a month for one whole year. Convert 1 year to months and multiply that to the $540 and you would get the total amount he pays in a year. Add the one time deposit payment and that would be the total amount he pays david jones back. Then it's simple subtraction to find the difference of the amount he had to pay vs the total cost of the furniture.

OpenStudy (anonymous):

wait i dont get it can u type the working out

OpenStudy (babalooo):

Which part confuses you? I'll start there.

OpenStudy (anonymous):

i dont get . Assuming the the 10% deposit is a one time payment, first calculate what the deposit amount is. 10% of 6000 is a tenth of 6000. What is that? Now In addition to that deposit, he pays $540 a month for one whole year. Convert 1 year to months and multiply that to the $540 and you would get the total amount he pays in a year. Add the one time deposit payment and that would be the total amount he pays david jones bac

OpenStudy (anonymous):

i just need a Working out

OpenStudy (babalooo):

I'm not here to do the work for you. Only here to simplify and clarify the question so you get it...

OpenStudy (anonymous):

is it 6000*10%=600 540 per month + 540 * 1/12=45 is that right

OpenStudy (babalooo):

The first part is right. The second part is not.

OpenStudy (babalooo):

He pays 540 a month for one whole year.

OpenStudy (anonymous):

so 540 * 1/12 = 45 right i accidenly added + sign instead = sign

OpenStudy (babalooo):

Now, we know one year = 12 months. And each month he pays $540. What is the total he pays in a year?

OpenStudy (babalooo):

why 1/12? The payment of $540 is a monthly payment. Each month, he pays $540.

OpenStudy (babalooo):

If you think about it.. if 600 and the 45 is correct, he won't be able to pay back the 6000 cost of the furniture right? I wouldn't sell you $6000 worth of furniture for $645

OpenStudy (anonymous):

12*540 is 6480 right 6480 -600 =5880

OpenStudy (babalooo):

why are you subtracting the $600 deposit? He also pays the $600 deposit. Unless it's refunded which is not included in the question. You should be finding the EXTRA amount he's paying.

OpenStudy (babalooo):

he pays deposit+ total amount (12*540) furniture cost = $6000